OpenStudy (anonymous):

Okay, so I've got this... y = (x^2 + 3) * sin pi.x^2 and z = (sin^2 pi.x)/(x^-2 + 3) for x = 0; 0:02; ... ; 10. Hence determine w = ((x^2 + 3) * sin pi.x^2 * sin^2 pi.x)/(x^-2 + 3) and plot w against x.... (I'm doing this on Matlab, so I'm not really sure how much help I can get for it, but my graph is coming out with just a horizontal line across the x-axis from 0-10).

OpenStudy (anonymous):

I didn't get a horizontal line. My graph looked like this.

OpenStudy (anonymous):

Here is my code: clear; clc; x = 0:0.02:10; y = (x.^2 + 3) .* sin(pi.*x.^2); z = sin(pi.*x).^2 ./ (x.^-2 + 3); w = y .* z; plot(x,w)

OpenStudy (anonymous):

Yeah, I got it in the end. I think I was missing one "." and that was it. Was getting so frustrated over it too. Thanks a lot